# Robustly optimal dynamics for active matter reservoir computing

### v1

@@ III.1 Around the critical damping regime: optimal performance
So far, it has been assumed that active matter reservoirs show optimal performance in a “critical” regime, around a fluid-to-gas-like phase transition [63], supporting the hypothesis of information processing being optimal around a phase transition [99]. Here we present a novel dynamical regime for improved reservoir computing with active matter: the critically damped regime. We uncover this regime by varying the parameters of the speed-controlling force $\mathbf{F}_{sc}$: the target agent speed $s$ and the strength $K_{sc}$ (see Eq. 4). Fig. 2a shows how well different active matter systems at different $(K_{sc},s)$ parameter combinations perform at predicting a chaotic input time series in a reservoir computing framework. We observe a broad region with high predictive performances which we demarcate with yellow dashed lines ($s=r(K_{sc})^{m}$ with power $m=1$ and coefficients $r_{AB}\approx 0.2$ and $r_{BC}\approx 70$). The optimal ratio $r$ that delivers performances around $P=0.884$ is located at $K_{sc}=10^{-5}$ and $s=5.0\cdot 10^{-5}$. This is a marked increase in performance from the previously reported optimal regime [63] (reproduced as $P=0.719$) indicated by the hexagon/“L” symbol in Fig. 2a. Videos of the driven active matter systems (see Appendix E, Tab. 3) reveal that, upon driver excitation, particles do not embark on long trajectories or oscillate across the simulation box before returning to their steady-state positions. Instead, the speed-controller setting ensures that agents swiftly return to the box center. This is why we call this regime “critically damped”. The quick return and the higher order in this regime can also be observed from the snapshots in Fig. 2b (square, triangle, circle symbols). The swarm maintains a coherent “bubble” shape with low mean speeds (see Fig. 19), even if perturbed by the driver.
@@
Below the lower yellow dashed line, agent dynamics begin to be dominated by the speed-controlling force. Here, the agents almost rest apart from local vibrational motion. For higher $K_{sc}$ and lower $s$ values (nabla symbol), agents perform alternating steps, where one direction (forward) is slightly stronger than the other (backwards), yielding an overall forward movement. Because agents do not interact strongly with each other or the driver, and instead oscillate alone, we call this regime “arrested”. The performance in Fig. 2a shows a sharp drop here, originating from the nearly complete absence of agent-driver interaction. Integrating the system with a smaller time step $\Delta t=0.002$ leads to smoother dynamics (see Supplementary Video Tab. 3/7) and removes the sharp performance drop for a local strip around the lower yellow boundary (see Fig. 16). Thus, the numerical integration time step plays a key role in this regime and its existence. Without a driver, both the underdamped and the critically damped regimes feature highly ordered steady states (see Tab. 7).
One simple, microscopic physical indicator of the difference between regimes is the mean squared displacement (MSD) after a delay time $\tau$. We choose as a suitable delay timescale the characteristic Lyapunov time of the driver, the Lorenz-63 system. Individual agent MSDs vary strongly even after multiple Lyapunov times in the underdamped regime, while MSDs remain low in the critically damped regime (see Fig. 14a). The agent-averaged MSDs presented in Fig. 2d are consistently higher in the underdamped regime compared with the critically damped regime. This low MSD region correlates with the high-performance prediction region in Fig. 2a. We also observe a stronger consistency of the types of responses visible in the critically damped regime in Supplementary Video Tab. 3/1-3, while this consistency is not clear in the underdamped regime in Supplementary Video Tab. 3/5. We hypothesize that underdamped system responses tend to generate more “different” spatio-temporal swarm patterns for similar driver trajectory patterns. This possibly impairs a key reservoir quality, the approximation property (similar external inputs should lead to similar reservoir states), which obfuscates the readout layer training. This could be one reason why the predictive performance is much higher in the critically damped regime than in the underdamped regime.
@@ III.2 Signatures from microscopic dynamics with and without driving
Figure 3: Characterizing intrinsic microscopic relaxation dynamics (a,b) as well as dynamical correlations under driving (c,d). Symbols correspond to parameter combinations changing speed-controller settings $(K_{sc},s)$ (and resulting damping dynamics) as shown in Fig. 2. (a,b) Intrinsic relaxation dynamics for a system without external driving in the absence of any collective effects (without agent-agent interactions, only subject to the homing force, the speed controller, and the sigmoid force wrapper described in Sect. II.1), the time step used is $\Delta t=0.02$. The system relaxes from an initial, uniformly random distribution to a steady state, analogous to a damped harmonic oscillator. (a) Structural relaxation of a population of non-interacting agents measured as the mean radial distance to the center of the simulation box, normalized by the initial radial distance at time $t_{0}=0$. (b) Corresponding dynamical relaxation: the relative deviation of agent speed from the set target agent speed $s$. Error bands indicate the standard error of the mean over $N=200$ agent trajectories. For reference, the Lyapunov timescale of the Lorenz-63 driver is shown in the time plots (a)-(c), i.e., $t^{\text{L63}}_{\text{lyap}}\approx 1.1$ [98] as a gray dashed line. Sustained oscillations around the center of the simulation box (underdamped regime, cross symbol) become suppressed when moving towards higher speed-controller strengths $K_{sc}$ and lower target agent speeds $s$ (diamond symbol). Critical damping is situated between the square and triangle symbols, at the threshold where a single oscillation disappears during the initial phase of the relaxation. (c,d) Dynamical correlations under driving for driven agents subject to the full active matter model described in Sect. II.1. (c) The absolute value of the velocity auto-correlations for selected damping dynamics. The optimal dynamics (around triangle) entail unique features described in the main text. (d) The connected velocity correlation function (CVC) for different parameter combinations. The optimally damped regime shows the strongest connected velocity fluctuation correlations and anti-correlations, measured spatially via a radial distance around each particle.
@@ III.2.1 Intrinsic damping dynamics: No driving
The underdamped, critically damped, and arrested dynamical regimes of the active matter substrates render visibly distinct RC performances. These terms are inspired by control theory [100] and by observing the damping behavior of individual agents (see Fig. 3). Intriguingly, one can identify these different regimes without the influence of an external driver, i.e., without an explicit injection of information. Even though we exclude the initial transient (“burn-in”) phase from training and inference in RC, it is highly informative about dynamics and associated computational properties: In Fig. 3a,b the relaxation behavior of a system without external driving is studied and categorized – in the absence of any collective effects (without agent-agent interactions, only subject to the homing force, the speed controller, and the sigmoid force wrapper described in Sect. II.1) (refer to Tab. 5 for the corresponding videos). These results show subtle features that are more visible than in the case of including interactions, displayed in the Appendix in Fig. 20. An ensemble of active agents relaxes from an initial uniformly random distribution to a steady-state configuration, similar to a damped harmonic oscillator. In sub-figure (a), structural relaxation is made visible via the mean radial distance to the center of the simulation box, normalized by the initial radial distance at $t_{0}=0.0$. Sustained oscillations around the center of the simulation box (underdamped regime, cross symbol) become suppressed when moving towards higher speed-controller strengths $K_{sc}$ and lower target agent speeds $s$, highlighted by the case in the diamond symbol.
@@
Fig. 3b shows the intrinsic microscopic relaxation dynamics as the relative deviation of agent speed from the set target agent speed from the initial condition. Seen in both (a) and (b), critical damping is situated between the square and triangle symbols, at the threshold where a single oscillation disappears during the initial phase of the relaxation, which becomes exponential for both the structure and dynamics at the square symbol. Further, the microscopic relaxation dynamics around the diamond symbol has an appearance of a power-law regime with superimposed oscillations. The oscillations decay and the power-law duration drastically shortens when introducing particle interactions Fig. 20d). In (b), the power-law(-like) regime is retained for the square and triangle symbols (nearly critically damped), and disappears for the circle symbol (overdamped). Small oscillations persist at later phases at the steady state. We refer to the crossover case near critical damping (triangle) as the “optimal damping” regime for active matter RC Fig. 2; it is the first overdamped parameter combination below critical damping.
@@
For the optimal damping case (triangle), there is a rapid relaxation to a low value in the structure (a), while in the dynamics, there is an intermediate stage exhibited that appears power-law-like. Intriguingly, for both cases of triangle and circle (more overdamped), the steady state values in dynamics are reached at the same time and with the same value. Yet, this is an effect of the lack of particle interactions: In the case of interacting particles, shown in the appendix in Fig. 20d,f, the optimally damped system (triangle) indeed converges earlier and more efficiently in dynamics than the case of the circle symbol. The combination of intrinsic exponential decay and a power law regime hence seems to have some utility for relaxing the dynamics (velocities) more effectively; it could, potentially, be a salient dynamical feature enabling favorable properties pertaining to the adaptability of the system, discussed later in Sect. IV. Refer to Tab. 6 for the corresponding videos for the interacting agents case.
@@
For the default time step of $\Delta t=0.02$ used here, agents at the nabla symbol parameter combination show arrested behavior. The mean normalized radial distances in (a) go above 1.0 because of periodic boundary conditions and the dynamical excitation in (b) attains a high value, an effect that disappears with a smaller timestep of $\Delta t=0.002$ (refer to Fig. 20a,b). In the latter, the onset of overdamping is shifted, which explains how the optimal computing regime is effectively widened in Fig. 16.
@@ III.2.2 Correlative dynamics under driving
We further characterize the system dynamics – now under the same external chaotic driving as in the RC setup and with full interactions described in Fig. 3c,d – in terms of spatial and temporal correlations of agent velocities, the velocity autocorrelation function (VAC), and the connected velocity correlation function (CVC). We take direct inspiration from the physics of metastable fluids and glassy systems, where one way of understanding dynamics is in the time-autocorrelation functions of particle mobilities [101, 102, 103, 104]. Velocities or related current autocorrelations render information on general transport and response properties of fluids [105, 106].
@@
The velocity autocorrelation (VAC) tells how a system’s microscopic dynamics, on average, relaxes; it has information on the second time derivative of basic structural scattering functions, i.e., the self-intermediate scattering function or the mean square displacements [107]; it is generally suitable for studying systems out of equilibrium. (It presumes, subtly, that long-time stationarity or quasi-repeatability of the dynamics persists, as it accumulates lag-time information over the absolute course of the forward evolution. Here, all data is collected after an initial “burn-in” period where the swarm system gets “warmed up” under driving.) Generally, the VAC can reveal signs of the heterogeneous dynamics visible in these systems, multiple processes responsible for relaxation, as well as signatures of coherent and incoherent motion, and provide a measure of the memory of a disturbance in a fluid [101, 102, 103, 104].
@@
In Fig. 3c, the absolute value of the velocity auto-correlations is plotted for selected variations of the speed-controller parameters (symbols) as shown in Fig. 2a. (Here, oscillations are turned into positive peaks with steep zero-crossings on the long-scale; the first peak always corresponds to a negative correlation, and positive peaks alternate thereafter.) The optimal dynamics (around case: triangle) entail initial short-time viscous or overdamped motion (an initial exponential decay). This is followed by harmonic-like motion with smooth, rounded peaks, exhibiting a frequency closely matching that of the external driving force (the time of the first zero crossing of the VAC). Intriguingly, a near-coincidence of the driver’s Lyapunov timescale, $t^{\text{L63}}_{\text{lyap}}\approx 1.1$ [98], with the minimum of the second (negative) peak.
@@
The strongly underdamped system (black cross) represents one (extreme) type of dynamics we encountered in Fig. 2a. What first strikes the eye is the slow decay of its envelope in the absolute VAC function (cross symbol), in Fig. 5a – a first agreement with our classification that the motion is underdamped, too, under chaotic driving. Moreover, it showcases harmonic-like motion with very long characteristic time scales: The successive peaks display a smooth, rounded profile, and the approximate wavelength (the distances between the sharp minima of the absolute VAC) is comparatively large, fitting more than half of a Lyapunov timescale indicated by the gray dotted line. As the first peak corresponds to a minimum in the VAC, partially coherent anti-correlated motion persists during the bulk of the characteristic chaotic timescale of the Lorenz-63 driver. There is a delay in the first zero crossing of the VAC – the first onset of anticorrelated velocities, which occurs later than for the other cases with more damping. The initial slope near $\tau=0$ is nearly flat and lacks a clear decay, indicating nearly ballistic or inertial-like motion at the shortest timescales–which can be visually confirmed as shown in Supplementary Video Tab. 3/5. The driven system in the underdamped regime exhibits slow expansive and contractive pulsations, which might generate a major part of these oscillations in the VAC (see the corresponding video). Further, the VAC for underdamped dynamics lacks a clear signature of separated dissipative (incoherent) dynamics: There is no sign of an exponential decay, i.e., viscous motion at short time-scales, which we do find for the overdamped dynamical regime. All these features together effectively can point to noticeable memory effects in the system, which would impair a key reservoir computing quality, the fading memory property. The system state should only depend on recent input signals and not on past inputs.
@@
Moving towards higher damping strengths, we observe much shorter timescales between extrema or zero crossings (i.e., shorter oscillations in the VAC). The optimally damped regime (triangle symbol), for example, displays an early first minimum and overall three minima in the absolute VAC within the Lyapunov time of the chaotic driver, compared to only two minima in the underdamped regime. Moreover, the peaks are nearly paraboloid-shaped for the former. Overall, around optimal damping, different kinds of dynamics are more discernible in delay time (i.e., short-time exponential relaxation and harmonic-like oscillations), and the time evolution is accompanied by strong mechanisms happening at shorter timescales than in the underdamped case. Note here that at the crossover (green circle) from the critically damped to the arrested regime, there are additional high-frequency oscillations in the VAC due to its vicinity to the arrested regime (nabla symbol, light green), where particles undergo arrested, back-and-forth oscillatory motion.
@@
Just above critical damping (at purple square), the absolute VAC begins to have overlapping features with that of the driver’s VAC (orange line), in particular, having the same first zero-crossing and mimicking the approximate frequency of the following two oscillations, to just before the Lyapunov time (there, the statistics are sensitive to sampling errors and a clear statement cannot be made). This appears to be an illustration of the system adapting dynamically to pronounced velocity fluctuations in the driver’s signal, possibly synchronizing some of its mechanisms to the driver [108].

### v2

@@ III.1 Around the critical damping regime: optimal performance
So far, it has been assumed that active matter reservoirs show optimal performance in a “critical” regime, around a fluid-to-gas-like phase transition [62], supporting the hypothesis of information processing being optimal around a phase transition [101]. Here we present a novel dynamical regime for improved reservoir computing with active matter: the critically damped regime. We uncover this regime by varying the parameters of the speed-controlling force $\mathbf{F}_{sc}$: the target agent speed $s$ and the strength $K_{sc}$ (see Eq. 4). Fig. 2a shows how well different active matter systems at different $(K_{sc},s)$ parameter combinations perform at predicting a chaotic input time series in a reservoir computing framework. We observe a broad region with high predictive performances which we define with yellow dashed lines ($s=r(K_{sc})^{m}$ with power $m=1$ and coefficients $r_{AB}\approx 0.2$ and $r_{BC}\approx 70$). The optimal ratio $r$ that delivers performances around $P=0.884$ is located at $K_{sc}=10^{-5}$ and $s=5.0\cdot 10^{-5}$. This is a significant increase in performance from the previously reported optimal regime [62] (reproduced as $P=0.719$) indicated by the hexagon/“L” symbol in Fig. 2a. Videos of the driven active matter systems (see Appendix E, Tab. 3) reveal that, upon driver excitation, particles do not embark on long trajectories or oscillate across the simulation box before returning to their steady-state positions. Instead, the speed-controller setting ensures that agents swiftly return to the box center. This is why we call this regime “critically damped”. The quick return and the higher order in this regime can also be observed from the snapshots in Fig. 2b (square, triangle, circle symbols). The swarm maintains a coherent droplet shape with low mean speeds (see Supplementary Fig. 19), even if perturbed by the driver.
@@
Below the lower yellow dashed line, agent dynamics begin to be dominated by the speed-controlling force. Here, the agents almost rest apart from local vibrational motion. For higher $K_{sc}$ and lower $s$ values (nabla symbol), agents perform alternating steps, where one direction (forward) is slightly stronger than the other (backwards), yielding an overall forward movement. Because agents do not interact strongly with each other or the driver, and instead oscillate alone, we call this regime “arrested”. The performance in Fig. 2a shows a sharp drop here, originating from the nearly complete absence of agent-driver interaction. Integrating the system with a smaller time step $\Delta t=0.002$ leads to smoother dynamics (see Supplementary Video Tab. 3/7) and removes the sharp performance drop for a local strip around the lower yellow boundary (see Supplementary Fig. 16). Thus, the numerical integration time step plays a key role in this regime and its existence. Without a driver, both the underdamped and the critically damped regimes feature highly ordered steady states (see Tab. 7).
One simple, microscopic physical indicator of the difference between regimes is the mean squared displacement (MSD) after a delay time $\tau$. We choose as a suitable delay timescale the characteristic Lyapunov time of the driver, the Lorenz-63 system. Individual agent MSDs vary strongly even after multiple Lyapunov times in the underdamped regime, while MSDs remain low in the critically damped regime (see Supplementary Fig. 14a). The agent-averaged MSDs presented in Fig. 2d are consistently higher in the underdamped regime compared with the critically damped regime. This low MSD region correlates with the high-performance prediction region in Fig. 2a. We also observe a stronger consistency of the types of responses visible in the critically damped regime in Supplementary Video Tab. 3/1-3, while this consistency is not clear in the underdamped regime in Supplementary Video Tab. 3/5. We hypothesize that underdamped system responses tend to generate more “different” spatio-temporal swarm patterns for similar driver trajectory patterns. This possibly impairs a key reservoir quality, the approximation property (similar external inputs should lead to similar reservoir states), which obfuscates the readout layer training. This could be one reason why the predictive performance is much higher in the critically damped regime than in the underdamped regime.
@@ III.2 Signatures from microscopic dynamics with and without driving
Figure 3: Characterizing intrinsic microscopic relaxation dynamics (a,b) as well as dynamical correlations under driving (c,d). Symbols correspond to parameter combinations changing speed-controller settings $(K_{sc},s)$ (and resulting damping dynamics) as shown in Fig. 2. (a,b) Intrinsic relaxation dynamics for a system without external driving in the absence of any collective effects (without agent-agent interactions, only subject to the homing force, the speed controller, and the sigmoid force wrapper described in Sect. II.1), the time step used is $\Delta t=0.02$. The system relaxes from an initial, uniformly random distribution to a steady state, analogous to a damped harmonic oscillator. (a) Structural relaxation of a population of non-interacting agents measured as the mean radial distance to the center of the simulation box, normalized by the initial radial distance at time $t_{0}=0$. Sustained oscillations around the center of the simulation box (underdamped regime, cross symbol) become suppressed when moving towards higher speed-controller strengths $K_{sc}$ and lower target agent speeds $s$ (diamond symbol). Critical damping is situated between the square and triangle symbols, at the threshold where a single oscillation disappears during the initial phase of the relaxation. (b) Corresponding dynamical relaxation: the relative deviation of agent speed from the set target agent speed $s$. Error bands indicate the standard error of the mean over $N=200$ agent trajectories. For reference, the Lyapunov timescale of the Lorenz-63 driver is shown in the time plots (a)-(c), i.e., $t^{\text{L63}}_{\text{lyap}}\approx 1.1$ [100] as a gray dashed line. (c,d) Dynamical correlations under driving for driven agents subject to the full active matter model described in Sect. II.1. (c) The absolute value of the velocity auto-correlations for selected damping dynamics. The optimal dynamics (around triangle) entail unique features described in the main text. (d) The connected velocity correlation function (CVC) for different parameter combinations. The optimally damped regime shows the strongest connected velocity fluctuation correlations and anti-correlations, measured spatially via a radial distance around each particle.
@@ III.2.1 Intrinsic damping dynamics: No driving
The underdamped, critically damped, and arrested dynamical regimes of the active matter substrates render visibly distinct RC performances. These terms are inspired by control theory [102] and by observing the damping behavior of individual agents (see Fig. 3). Intriguingly, one can identify these different regimes without the influence of an external driver, i.e., without an explicit injection of information. Even though we exclude the initial transient (“burn-in”) phase from training and inference in RC, it is highly informative about dynamics and associated computational properties: In Fig. 3a,b the relaxation behavior of a system without external driving is studied and categorized – in the absence of any collective effects (without agent-agent interactions, only subject to the homing force, the speed controller, and the sigmoid force wrapper described in Sect. II.1) (refer to Tab. 5 for the corresponding videos). These results show subtle features that are more visible than in the case of including interactions, displayed in the Supplementary Fig. 20. An ensemble of active agents relaxes from an initial uniformly random distribution to a steady-state configuration, similar to a damped harmonic oscillator. In sub-figure (a), structural relaxation is made visible via the mean radial distance to the center of the simulation box, normalized by the initial radial distance at $t_{0}=0.0$. Sustained oscillations around the center of the simulation box (underdamped regime, cross symbol) become suppressed when moving towards higher speed-controller strengths $K_{sc}$ and lower target agent speeds $s$, highlighted by the diamond symbol.
@@
Fig. 3b shows the intrinsic microscopic relaxation dynamics as the relative deviation of agent speed from the set target agent speed from the initial condition. Seen in both (a) and (b), critical damping is situated between the square and triangle symbols, at the threshold where a single oscillation disappears during the initial phase of the relaxation, which becomes exponential for both the structure and dynamics at the square symbol. Further, the microscopic relaxation dynamics around the diamond symbol has an appearance of a power-law regime with superimposed oscillations. The oscillations decay and the power-law duration drastically shortens when introducing particle interactions Fig. 20d). In (b), the power-law(-like) regime is retained for the square and triangle symbols (nearly critically damped), and disappears for the circle symbol (overdamped). Small oscillations persist at later phases at the steady state. We refer to the crossover case near critical damping (triangle symbol) as the “optimal damping” regime for active matter RC Fig. 2; it is the first overdamped parameter combination below critical damping.
@@
For the optimal damping case (triangle symbol), there is a rapid relaxation to a low value in the structure (a), while in the dynamics, there is an intermediate stage exhibited that appears power-law-like. Intriguingly, for both cases of triangle and circle (more overdamped), the steady state values in dynamics are reached at the same time and with the same value. Yet, this is an effect of the lack of particle interactions: In the case of interacting particles, shown in the Supplementary Fig. 20d,f, the optimally damped system (triangle) indeed converges earlier and more efficiently in dynamics than the case of the circle symbol. The combination of intrinsic exponential decay and a power-law-like regime, hence, seems to have some utility for relaxing the dynamics (velocities) more effectively. It could be a dynamical feature enabling favorable properties of the system, discussed later in Sect. IV. Refer to Tab. 6 for the corresponding videos for the interacting agents case.
@@
For the default time step of $\Delta t=0.02$ used here, agents at the nabla symbol parameter combination show arrested behavior. The mean normalized radial distances in (a) go above 1.0 because of periodic boundary conditions and the dynamical excitation in (b) attains a high value, an effect that disappears with a smaller timestep of $\Delta t=0.002$ (refer to Fig. 20a,b). In the latter, the onset of overdamping is shifted, which could explain how the optimal computing regime is effectively widened in Fig. 16.
@@ III.2.2 Correlative dynamics under driving
We further characterize the system dynamics – now under the same external chaotic driving as in the RC setup and with full interactions described in Fig. 3c,d – in terms of spatial and temporal correlations of agent velocities, the velocity autocorrelation function (VAC), and the connected velocity correlation function (CVC). We take direct inspiration from the physics of metastable fluids and glassy systems, where one way of understanding dynamics is the time-autocorrelation functions of particle mobilities [103, 104, 105, 106]. Velocities or related current autocorrelations render information on general transport and response properties of fluids [107, 108].
@@
The velocity autocorrelation (VAC) tells how a system’s microscopic dynamics, on average, relaxes; it has information on the second time derivative of basic structural scattering functions, i.e., the self-intermediate scattering function or the mean square displacements [109]. It is generally suitable for studying systems out of equilibrium. It requires that long-time stationarity or quasi-repeatability of the dynamics persists, as it accumulates lag-time information over the absolute course of the forward evolution. Here, all data is collected after an initial “burn-in” period where the swarm system gets “warmed up” under driving. Generally, the VAC can reveal signs of the heterogeneous dynamics visible in these systems, multiple processes responsible for relaxation, as well as signatures of coherent and incoherent motion, and provide a measure of the memory of a disturbance in a fluid [103, 104, 105, 106].
@@
In Fig. 3c, the absolute value of the velocity auto-correlations is plotted for selected variations of the speed-controller parameters (symbols) as shown in Fig. 2a. (Here, oscillations are turned into positive peaks with steep zero-crossings on the long-scale; the first peak always corresponds to a negative correlation, and positive peaks alternate thereafter.) The optimal dynamics (around the triangle symbol) entail initial short-time viscous or overdamped motion (an initial exponential decay). This is followed by harmonic-like signals with rounded peaks, exhibiting a shorter frequency closely matching that of the external driving force (the time of the first zero crossing of the VAC). Intriguingly, a near-coincidence of the driver’s Lyapunov timescale, $t^{\text{L63}}_{\text{lyap}}\approx 1.1$ [100], with the minimum of the second (negative) peak.
@@
The strongly underdamped system (black cross) represents one (extreme) type of dynamics we encountered in Fig. 2a. What first strikes the eye is the slow decay of its envelope in the absolute VAC function (cross symbol), in Fig. 5a – a first agreement with our classification that the motion is underdamped, too, under chaotic driving. Moreover, it showcases harmonic-like motion with very long characteristic time scales: The successive peaks display a rounded profile, and the approximate wavelength (the distances between the sharp minima of the absolute VAC) is comparatively large, fitting more than half of a Lyapunov timescale indicated by the gray dotted line. As the first peak corresponds to a minimum in the VAC, partially coherent anti-correlated motion persists during the bulk of the characteristic chaotic timescale of the Lorenz-63 driver. There is a delay in the first zero crossing of the VAC – the first onset of anticorrelated velocities, which occurs later than for the other cases with more damping. The initial slope near $\tau=0$ is nearly flat and lacks a clear decay, indicating nearly ballistic or inertial-like motion at the shortest timescales–which can be visually confirmed as shown in Supplementary Video Tab. 3/5. The driven system in the underdamped regime exhibits slow expansive and contractive pulsations, which might generate a major part of these oscillations in the VAC (see the corresponding video). Further, the VAC for underdamped dynamics lacks a clear signature of separated dissipative (incoherent) dynamics: There is no sign of an exponential decay, i.e., viscous motion at short time-scales, which we do find for the overdamped dynamical regime. All these features together effectively can point to noticeable memory effects in the system, which would impair a key reservoir computing quality, the fading memory property. The system state should only depend on recent input signals and not on past inputs.
@@
Moving towards higher damping strengths, we observe much shorter timescales between extrema or zero crossings (i.e., shorter oscillations in the VAC). The optimally damped regime (triangle symbol), for example, displays an early first minimum and overall three minima in the absolute VAC within the Lyapunov time of the chaotic driver, compared to only two minima in the underdamped regime. Moreover, the peaks are nearly paraboloid-shaped for the former. Overall, around optimal damping, a diversity of discernible types of dynamics appears (i.e., short-time exponential relaxation and harmonic-like oscillations). Note here that at the crossover (green circle) from the critically damped to the arrested regime, there are additional high-frequency oscillations in the VAC due to its vicinity to the arrested regime (nabla symbol, light green), where particles undergo arrested, back-and-forth oscillatory motion.
@@
Just above critical damping (at purple square), the absolute VAC begins to have overlapping features with that of the driver’s VAC (orange line), in particular, having the same first zero-crossing and mimicking the approximate frequency of the following two oscillations, to just before the Lyapunov time (there, the statistics are sensitive to sampling errors and a clear statement cannot be made). The physical system may be synchronizing some of its mechanisms to the driver [110].